OpenStudy (anonymous):

2(6t+1) - 4(t-1) = 0 Please show work out.

OpenStudy (anonymous):

t = -3/4

OpenStudy (anonymous):

2(6t+1)-4(t-1)=0: (distribute) 12t+2-4t+4=0 (combine like terms) 8t+6=0 (move the "6") 8t=-6 (divide both sides by 8) t = -3/4